Sunday 03/16/14

10:30a – 6:00p

Goals

  • flip fly stocks (done)
  • passage cells (done)
  • scheduling
    • christia (May 17-20)
    • dad’s conference (April 25th)
    • dentist (May 7th)

More data processing!

  • running splitdax on 140310_L4E4 data
  • running bead averaging on 140310_L4E4 data
  • analyzing averaged beads from 140310_L4E4
  • Not all 140310_L4E4 data copied from STORM4! copying remaining data. Need to relaunch all analysis 🙁
  • running splitdax on dual color data — currently saves the 647 channel from the 750 movies because we haven’t told it to distinguish. This makes all the running a bit slower.

Issues with data

  • 2014-01-14_F03-6_F04-7 beads alignment is clearly off — systematic shift in 750 647 observed
  • 2014-01-15_F04-6_F03-7. bead alignment is okay, substantial shift (8-15 um?) in acquiring conventional images to STORM. Can’t use conventional masks! 🙁
  • 2014-01-17_F03-6_F05-7. 750 data not analyzed (only beads are analyzed in 750, hopefully using the correct parameters).

Project 2

  • looking at beads and setting up data run with Hao

Ph Project

  • Ajaz found bug in k_d equation. Rerunning simulations. I think this actually expands the parameter space that gives the decreasing Ph behavior (as suits the logic). Previous dissociation rate had wrong dependence on mutant PH concentration in the cluster.
  • more explorations.
This entry was posted in Summaries and tagged , , . Bookmark the permalink.